OpenStudy (anonymous):

calculate the work required to be done to stop a car of 1500 kgg moving at a velocity of 60km/h.

OpenStudy (anonymous):

m= 1500kg u=60km/h = 60×1000m/3600s = 50/3m/s work required to be done to stop the car W = change in K.E. of the car =1/2mv² - 1/2m (0)² = 1/2mu² putting d value v get W= 1/2×1500× [50/3]² = 208333.3 J